OpenStudy (anonymous):

Solve 4(x + 1) + 8 = 24. 3.75 -2 5 3

OpenStudy (anonymous):

3

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Thanks can you guys help me with a few others? :)

OpenStudy (anonymous):

Solve 3c + 4 = -20. -5.3 6 8 -8

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

3c+4=-20 3c=-20-4 3c=-24 c=-8